### v2

# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

Abstract
The use of copyrighted books for training AI models has led to numerous lawsuits from authors concerned about AI’s ability to generate derivative content. Yet it’s unclear whether these models can generate high quality literary text while emulating authors’ styles/voices. To answer this we conducted a preregistered study comparing MFA-trained expert writers with three frontier AI models: ChatGPT, Claude, and Gemini in writing up to 450 word excerpts emulating 50 award-winning authors’ (including Nobel laureates, Booker Prize winners, and young emerging National Book Award finalists) diverse styles. In blind pairwise evaluations by 159 representative expert (MFA candidates from top U.S. writing programs) and lay readers (recruited via Prolific), AI-generated text from in-context prompting was strongly disfavored by experts for both stylistic fidelity (odds ratio [OR]=0.16, $p<10^{-8}$) and writing quality (OR=0.13, $p<10^{-7}$) but showed mixed results with lay readers. However, fine-tuning ChatGPT on individual author’s complete works completely reversed these findings: experts now favored AI-generated text for stylistic fidelity (OR=8.16, $p<10^{-13}$) and writing quality (OR=1.87, p=0.010), with lay readers showing similar shifts. These effects are robust under cluster-robust inference and generalize across authors and styles in author-level heterogeneity analyses. The fine-tuned outputs were rarely flagged as AI-generated (3% rate versus 97% for in-context prompting) by state-of-the-art AI detectors. Mediation analysis reveals this reversal occurs because fine-tuning eliminates detectable AI stylistic quirks (e.g., cliché density) that penalize in-context outputs, altering the relationship between AI detectability and reader preference. While we do not account for additional costs of human effort required to transform raw AI output into cohesive, publishable prose, the median fine-tuning and inference cost of \$81 per author represents a dramatic 99.7% reduction compared to typical professional writer compensation. Author-specific fine-tuning thus enables non-verbatim AI writing that readers prefer to expert human writing, thereby providing empirical evidence directly relevant to copyright’s fourth fair-use factor, the “effect upon the potential market or value” of the source works.
Keywords: Generative AI, Copyright Law, Fair Use, Future of Work, AI Detection, AI and Society, Behavioral Science, Labor Market Impact
1 Introduction
The U.S. publishing industry supports hundreds of thousands of jobs while generating \$30 billion in yearly revenue, contributing to the larger American copyright sectors that account for \$2.09 trillion in annual GDP contributions (?). Adult Fiction and nonfiction books alone accounted for \$6.14 billion in 2024 (?). This economically important sector now faces an unprecedented challenge: its core products have become essential training data for generative-AI systems. Models trained on well-edited books produce more coherent, accurate responses—something crucial to creating the illusion of intelligence (?). Most technology companies building AI use massive datasets of books, typically without permission or licensing (?), and frequently from illegal sources. In the recent copyright lawsuit Bartz v. Anthropic (?), Judge Alsup noted that Anthropic acquired at least five million books from LibGen and two million from Pirate Library Mirror (PiLiMi). Anthropic also used Books3111Now removed after a legal complaint by anti-piracy group, the Rights Alliance.—a dataset of approximately 191,000 books also used by Meta and Bloomberg to train their language models. [^2] The Bartz v. Anthropic lawsuit also revealed how Anthropic cut millions of print books from their bindings, scanned them into digital files, and threw away the originals solely for the purpose of training Claude. This unauthorized use has sparked outrage among authors (?), triggering dozens of lawsuits against technology companies including OpenAI, Anthropic, Microsoft, Google, and Meta.
Generative AI systems such as ChatGPT that can be prompted to create new text at scale are qualitatively unlike most historical examples of automation technologies (?). They can now solve Olympiad-level Geometry (?), achieve an impressive rating of 2700 on Codeforces, one of the most challenging coding competition platforms (?), and deliver medical guidance that meets professional healthcare standards (?, ?). Recent findings from Microsoft’s Occupational Implications of Generative AI (?), Anthropic’s Economic Index (?) and OpenAI (?) reveal AI usage primarily concentrates in writing tasks. This concentration threatens creative writing professionals in particular—novelists, poets, screenwriters, and content creators who shape cultural narratives and human expression. Based on U.S. Bureau of Labor Statistics May 2023 national estimates, creative writing constitutes almost 50% of writing jobs (?), making these positions especially vulnerable to GenAI-based automation, as the writing community has warned (?).
While it is widely established that most frontier large language models (LLMs) have been trained on copyrighted books, it remains unclear whether such training can produce expert-level creative writing. Past research has shown that AI cannot produce highbrow literary fiction or creative nonfiction through prompting alone when compared to professionally trained writers (?). More recent work from (?) demonstrates that AI-generated creative writing still remains characterized by clichés, purple prose, and unnecessary exposition. Additionally, relying on GenAI for creative writing reduces the collective diversity of novel content (?). AI often produces formulaic, mediocre creative writing because it lacks the distinctive personal voice that typically distinguishes one author from another (?). As Pulitzer Fiction finalist Vauhini Vara observes, “ChatGPT’s voice is polite, predictable, inoffensive, upbeat. Great characters, on the other hand, aren’t polite; great plots aren’t predictable; great style isn’t inoffensive; and great endings aren’t upbeat” (?). To address this limitation, practitioners now increasingly prompt AI systems to perform style/voice mimicry by emulating specific writers’ choices (?). This practice has become so common that a fantasy author recently published a novel containing an accidentally included AI prompt requesting emulation of another writer’s style (?). While the effectiveness of such stylistic emulation remains contested, the more pressing question concerns whether style/voice mimicry genuinely improves AI-generated text quality and whether readers—both experts and non-experts—perceive these improvements as meaningful.
To address this question, we conducted a preregistered behavioral study comparing MFA-trained expert writers with froniter large language models. Historically, top MFA programs have produced many prizewinning American writers  (?). Eminent literary agent Gail Hochman of Brandt & Hochman said, “We look favorably on anyone who has an M.F.A., simply because it shows they’re serious about their writing”  (?). This elite sample of MFA trained expert writers provides a conservative test—if AI can compete with the best emerging talent, the disruption to average writers is likely even greater. We selected closed-source LLMs readily accessible to users without technical expertise: GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. [^3] We also tried open-weight Llama 3.1 model but its empirical performance was not good at following long context instructions at the time of our study. Additionally our results would hold across different LLM versions in the future as LLM outputs have been shown to be homogeneous and have not improved over time (?, ?, ?, ?). Both human experts and LLMs were given the same task: write an excerpt of up to 450 words emulating the style and voice of one of 50 internationally acclaimed authors, including Nobel laureates, Booker Prize winners, and Pulitzer Prize winners, spanning multiple continents and cultures. [^4] Among others, our study includes Nobel laureates Han Kang and Annie Ernaux; Booker Prize winners Salman Rushdie, Margaret Atwood, and George Saunders; and Pulitzer Prize winners Junot Diaz and Marilynne Robinson. We tested two AI conditions: (1) in-context prompting, where models received the same instructions as human experts, and (2) fine-tuning, where models were additionally trained on each author’s complete oeuvre. [^5] For authors writing in non-English languages (Han Kang, Yoko Ogawa, Annie Ernaux, Haruki Murakami), we used the same translator’s work across all books to maintain voice consistency. Expert and lay readers performed blind pairwise evaluations (?, ?, ?) of <<Human-AI>> excerpts on writing quality and stylistic fidelity. This design addresses three preregistered research questions: (1) Can AI match expert performance in writing quality and stylistic fidelity across both conditions? (2) Do expert and lay readers show similar preference patterns? (3) Does AI detectability correlate with human quality judgments, and does fine-tuning remove this correlation? Our full experimental setup is shown in Figure 1.
Figure 1: Figure showing our study design. (1) Select a target author and prompt. (2) Generate upto 450-word candidate excerpts from MFA experts and from LLMs under two settings: in-context prompting (instructions + few-shot examples) and author-specific fine-tuning (model fine-tuned on that author’s works). (3) Readers (experts and lay) perform blinded, pairwise forced-choice evaluations on two outcomes: stylistic fidelity to the target author and overall writing quality. Pair order and left/right placement are randomized on every trial.
2 Results
Figure 2: (A-B) Forest plots showing odds ratios (OR) and 95% confidence intervals comparing AI and human experts in pairwise evaluations of stylistic fidelity (A) and writing quality (B) where values >>1 favor AI and values >>1 favor humans. Expert readers show preference for human writing when prompted in an in-context setting (OR = 0.16 and 0.13) but that changes when AI is fine-tuned (OR = 8.16 and 1.87). Lay readers have a harder time discriminating, given how they prefer the quality of AI writing even for in-context prompting (OR = 1.55). (C-D) Probability of choosing AI excerpts across individual language models for stylistic fidelity (C) and writing quality (D). Error bars represent 95% confidence intervals. Dashed line indicates chance performance (50%). (E) AI detection accuracy with chosen threshold of $\tau$=0.9 using two state-of-the-art AI detectors (Pangram and GPTZero). Human written text was never misclassified (0.00), in-context AI was detected with 97% accuracy by Pangram and 91% by GPTZero, but fine-tuned AI evaded detection 97% of the time (0.03) for Pangram and 100% of the time in case of GPTZero. (F) Relationship between AI detectability (Pangram) and preference for writing quality across detection score bins. For in-context prompting setup, higher detection scores correlated with lower AI preference (negative slope). This relationship disappeared when AI is fine-tuned (flat slopes). Fine-tuning on an author’s complete oeuvre eliminates stylometric “AI” quirks while achieving expert-level performance. n = 28 expert readers, 131 lay readers; 3,840 pairwise comparisons with robust clustered standard errors.
2.1 Overall Performance Comparisons
Our final data consist of 3,840 paired-choice tasks, with judgments from 28 MFA experts and 131 lay readers. We fit a logit model for each outcome and condition and include dummies for writer type, reader group, and their interaction. We further employ CR2 cluster-robust standard errors (?) clustered at the reader-level to account for within-reader correlation in ratings. Our hypotheses, outcomes, design, and analysis closely follow our OSF pre-registration (SI Sections S4-S8); deviations are detailed in SI Section S9.
Figure 2A–B presents the odds ratios, with corresponding predicted probabilities shown in Figure 2C–D. Under in-context prompting, expert readers demonstrated strong preference for human-written text. Odds ratios were 0.16 (95% CI: 0.08–0.29, $p<10^{-8}$) for stylistic fidelity and 0.13 (95% CI: 0.06–0.28, $p<10^{-7}$) for writing quality, indicating six- to eight-fold preferences for human excerpts (Fig. 2A–B). Lay readers showed no significant preference regarding stylistic fidelity (OR = 0.86, 95% CI: 0.62–1.19, $p=0.37$) but favored AI-generated text for writing quality (OR = 1.55, 95% CI: 1.09–2.21, $p=0.014$), selecting AI excerpts in 61% of writing quality trials (Fig. 2C–D). Inter-rater agreement reflected this divergence: expert readers achieved $\kappa=0.58$ for stylistic fidelity and $\kappa=0.41$ for writing quality, while lay readers showed minimal agreement among themselves ($\kappa=0.12$ and $\kappa=0.15$, respectively). [^6] Given that lay assessments of literary quality and style reflect inherently diverse tastes, we anticipated lower inter-rater agreement compared to expert readers. The writer-type $\times$ reader-type interaction was significant for both outcomes ($\chi^{2}_{(3)}=24.9$, $p=1.6\times 10^{-5}$ for fidelity; $\chi^{2}_{(3)}=37.6$, $p=3.5\times 10^{-8}$ for quality).
Fine-tuning on authors’ complete works reversed these preferences. For expert readers, the odds of selecting the AI excerpt were 8.16 times the odds of selecting the human excerpt for stylistic fidelity (OR =8.16=8.16, 95% CI: 4.69–14.2, $p<10^{-13}$) and 1.87 times the odds for writing quality (OR =1.87=1.87, 95% CI: 1.16–3.02, $p=0.010$). Lay readers showed comparable shifts in their preferences (stylistic fidelity OR =8.29=8.29, 95% CI: 5.15–13.3, $p<10^{-17}$; writing quality OR =2.42=2.42, 95% CI: 1.64–3.57, $p<10^{-5}$). Model-based predicted AI win probabilities converged across groups to about 0.74 for stylistic fidelity and 0.58–0.61 for writing quality (Fig. 2C–D), and the writer-type $\times$ reader-type interaction was no longer significant in the fine-tuned models (both $\chi^{2}_{(1)}<1$, $p>0.40$). Inter-rater agreement among experts increased ($\kappa=0.67$ for writing quality; $\kappa=0.54$ for stylistic fidelity), while agreement among lay readers remained low ($\kappa=0.07$ and $\kappa=0.22$, respectively).
[…]
2.2 AI Detection and Stylometric Analysis
We probe whether differences in AI detectability can account for these preference reversals. Pangram, a state-of-the-art AI detection tool (?, ?, ?), correctly classified 97% of in-context prompted texts as machine-generated but only 3% of fine-tuned texts were classified as AI-generated (Fig. 2E). [^7] GPTZero, another state-of-the-art AI detection tool showed comparable performance but had higher false-positive-rate, so for our subsequent analyses we stuck to Pangram.
Higher AI-detection scores strongly predicted lower preference rates among expert readers in the in-context prompting condition. For stylistic fidelity, each unit increase in detection score reduced the odds of selecting an excerpt by a factor of 6.3 ($\beta=-1.85\pm 0.29$, $p<10^{-9}$); a similar pattern held for writing quality ($\beta=-2.01\pm 0.33$, $p<10^{-9}$). Fine-tuning largely eliminated this negative relationship between detectability and preference (Pangram $\times$ Setting for style: $\beta=2.56\pm 0.81$, $p=0.002$; for quality: $\beta=2.90\pm 0.88$, $p<0.001$). Two-stage mediation analysis (Fig. 4A) demonstrated that stylometric features, particularly cliché density (See Section S3.2 in SI), mediated 16.4% of the detection effect on preference before fine-tuning but a statistically insignificant $1.3\%$ (95% CI includes zero) afterward, indicating that fine-tuning eliminates rather than masks artificial stylistic signatures.
2.3 Author-Level Performance Heterogeneity
Next, we disaggregated our data to unpack heterogeneity at the level of individual authors. Of 30 fine-tuned author models, 27 exceeded parity for stylistic fidelity (median win rate = 0.74, IQR: 0.63–0.86) and 23 for writing quality (median = 0.58, IQR: 0.54–0.74). Using 95% Jeffreys intervals, fine-tuned models significantly outperformed human writers for 19 authors on stylistic fidelity and 10 authors on writing quality (Fig. 3A). These performance differences showed no systematic relationship with fine-tuning corpus size (Fig. 3B). The fine-tuning premium, i.e., the increase in AI preference rates relative to in-context prompting, ranged from 13.7-13.7 to +70.8+70.8 percentage points for stylistic fidelity (29 of 30 positive) and from 20.8-20.8 to +50.0+50.0 percentage points for writing quality (22 of 30 positive). This “premium” likewise showed no correlation with fine-tuning corpus size (Pearson $r<0.1$ for both outcomes; Fig. 4B).
Figure 4: Fine-tuning substantially reduces stylometric signatures of AI text, improves stylistic fidelity and perceived writing quality over in-context prompting, and substantially cuts costs of producing first draft versus professional writers. (A) Mediation analysis linking stylometric features $\rightarrow$ AI-detector score $\rightarrow$ human preference. Standardized logistic coefficients with 95% CIs are shown for three features for in-context prompting (red) and fine-tuned models (green). Cliché density mediates 16.4% of the detector effect on choice for in-context prompting but only 1.3% for author fine-tuned models; all three features together mediate 25.4% vs -3.2% for in-context prompted and fine-tuned models respectively. (B) “Fine-tuning premium,” defined as $\Delta$ = P(prefer fine-tuned over human) - P(prefer in-context over human), as a function of fine-tuning corpus size. Top: stylistic fidelity; bottom: writing quality. Points are authors; colors denote improvement (green, $\Delta$¿0), no change (gray), or degradation (red, $\Delta<0$). Median $\Delta$: +41.7 (fidelity) and +16.7 percentage points (quality). (C) Cost to produce 100,000 words of raw text vs. publishable prose. Expert writers in our study would earn \$25,000 for a 100k-word novel-length manuscript (red). By contrast, AI pipelines can generate 100k words of raw text for \$25–\$276 depending on fine-tuning corpus size (green bars = fine-tuning; hatched = in-context prompting, \$3). This figure reflects direct compute/API costs only, not the additional human steering, chunking, and editing required to turn raw AI text into a cohesive publishable work. Authors ordered by total AI cost.
2.4 Cost Analysis
The weak dependence of performance on fine-tuning corpus size has direct economic implications. Model fine-tuning and inference costs ranged from \$25 to \$276 per author (median = \$81), assuming API-based fine-tuning at \$25 per million tokens plus \$3 for generating 100,000 words of raw text (Fig. 4C). These costs represent approximately 0.3% of what expert writers in our study would charge for a novel-length (100,000 words) manuscript. It should be noted that this comparison reflects raw generation costs before the human steering and editing required to transform AI outputs into publishable works. Despite this caveat, the minimal investment yielded outputs that achieved expert-level performance for most authors. Performance gains were uncorrelated with both corpus size and fine-tuning cost, indicating that computational scale did not drive improvements. The 99.7% reduction in raw generation costs, coupled with superior quality ratings for the majority of authors, underscores the potential for substantial producer surplus shifts and market displacement.
3 Discussion
What are the implications of this research for the copyright infringement claims that authors have brought against AI companies alleging unauthorized use of their books in training datasets (?, ?, ?)? These cases raise the question whether copying millions of copyrighted books for AI training constitutes fair use when the resulting outputs do not themselves reproduce the copied works. The most significant consideration in evaluating this question is the fourth fair use factor: “the impact of the use upon the actual or potential market for the copyrighted work.” [^8] The fair use provision (section 107) of the US Copyright Act directs courts to consider four factors: 1. the purpose and character of the use, including whether such use is of a commercial nature or is for nonprofit educational 2. the nature of the copyrighted work; 3. the amount and substantiality of the portion used in relation to the copyrighted work as a whole; and 4. the effect of the use upon the potential market for or value of the copyrighted work. Courts understand this factor to concern the extent to which the works produced by the copying substitute for the author’s work.
Courts understand this factor to concern the extent to which works produced through copying serve as market substitutes for the original author’s work (?, ?). Our study has shown that AI-generated excerpts from in-context prompted models pre-trained on vast internet corpora (including millions of copyrighted works) were strongly disfavored by expert readers for both quality and stylistic fidelity, though lay readers showed no clear preference. By contrast, when models are fine-tuned on curated datasets—consisting solely of individual author’s complete works—both experts and lay readers decisively preferred the AI-generated excerpts over human-written examples. Moreover, the fine-tuned excerpts proved almost undetectable as AI-generated text (particularly when compared to outputs from in-context prompted models), while consistently surpassing human writing in blind evaluations.
At first glance, a legal analyst might conclude that our findings are irrelevant to fair use because the outputs from the blind pairwise evaluation do not reproduce the copied works. While they may exhibit comparable literary quality and high stylistic fidelity to the originals, copyright law does not protect authors’ style—only their expression (?, ?). These outputs may offer credible substitutes for an author’s works, but so do human-authored works inspired by prior works. However, there is an important difference between human and AI-generated emulations: humans read; AI systems copy. Unlike human memory, which is not a verbatim storage device, all AI-generation requires predicate copying despite rhetoric equating human learning with machine “learning.” (?, ?, ?)
The Copyright Office has recognized that such predicate copying may cause cognizable market harm through competing works that the inputs enable, potentially flooding the market and causing “market dilution” (?). While acknowledging this “market dilution” approach to the fourth fair use factor as “uncharted territory,” the Office determined that both the statutory language and underlying concerns of the Copyright Act warrant this inquiry.999“…The speed and scale at which AI systems generate content pose a serious risk of diluting markets for works of the same kind as in their training data. That means more competition for sales of an author’s works and more difficulty for audiences in finding them. If thousands of AI-generated romance novels are put on the market, fewer of the human-authored romance novels that the AI was trained on are likely to be sold…Market harm can also stem from AI models’ generation of material stylistically similar to works in their training data…” The Office emphasized that the effect of the copying impacts extant works by putting them in competition with AI-generated outputs. Copyrighted works become fodder for new productions targeting the same markets. Crucially, the Office did not claim that competing AI outputs copy the inputted works; rather, it examined the economic consequences of the predicate copying that enables these competing outputs. This focus on inputs remains essential because, absent the initial copying, no infringement action exists against flooding markets with independently generated works—if a thousand humans write romance novels after reading Barbara Cartland’s novels, they compete but do not infringe. The Copyright Office’s expansive interpretation of “potential market for or value of the copied work” suggests that fair use might not excuse predicate copying even when it doesn’t show up in the end product, if the copying’s effect substitutes for source works.
In Kadrey v. Meta (?), an infringement action brought by 13 book authors against the copying of their books into the database underpinning Meta’s Llama LLM, Judge Chhabria granted summary judgment to Meta but accepted the theory of market dilution:101010[I]ndirect substitution is still substitution: If someone bought a romance novel written by an LLM instead of a romance novel written by a human author, the LLM-generated novel is substituting for the human-written one…This case involves a technology that can generate literally millions of secondary works, with a miniscule fraction of the time and creativity used to create the original works it was trained on. No other use…has anything near the potential to flood the market with competing works the way that LLM training does. Judge Chhabria effectively provided a road map for what authors would have to show to persuade a court that the AI inputs diluted their markets: First, is the AI system “capable of generating” substitutional books? Second, what are the markets for the plaintiffs’ books, and do the AI-generated books compete in those markets? “Third, what impact does this competition actually have on sales of the books it competes with?…Whatever the effects have been thus far, are they likely to increase in the future, as more and more AI-generated books are written, and as LLMs get better and better at writing human-like text? Fourth, how does the threat to the market for the plaintiffs’ books in a world where LLM developers can copy those books compare to the threat to the market for the plaintiffs’ books in a world where the developers can’t copy them?”
Our study’s findings concerning reader preferences between human-authored and AI-generated works bear on all four considerations. They also demonstrate how LLMs have already gotten “better and better at writing human-like text.” While Judge Chhabria speculated that the distinctiveness of an author’s style renders works by well-known authors less susceptible to substitution, [^11] Judge Chhabria observed that market dilution would vary by author prominence: established authors with dedicated readerships (like Agatha Christie) would likely face minimal substitution, while AI-generated books could crowd out lesser-known or emerging authors, potentially preventing “the next Agatha Christie from getting noticed or selling enough books to keep writing.” our work suggests otherwise. If readers in fact prefer AI-generated emulations of authors whose market value lies in their distinctive voices, then the prospect of competition, especially from outputs of fine-tuned datasets, appears to be considerable. The comparatively low production costs of AI-generated texts relative to paying human authors (as shown in Figure 4C) further enhances the likelihood that AI platforms will in fact dilute the market for human-authored work.
These findings suggest that the creation of fine-tuned LLMs consisting of the collected copyrighted works (or a substantial number) of individual authors should not be fair use if the LLM is used to create outputs that emulate the author’s works. As the Copyright Office observed, “[f]ine-tuning…usually narrows down the model’s capabilities and might be more aligned with the original purpose of the copyrighted material,” (?) and thus both less “transformative,” and more likely to substitute for it. By contrast, the LLMs employed for in-context prompting do not target particular authors, and therefore can be put to a great variety of uses that do not risk diluting those authors’ markets. Their claim to fair use seems accordingly stronger. But those models can generate author-emulations, and our study has shown that at least as to lay readers, those outputs can substitute.
A reasonable solution might allow the inclusion of copyrighted works in the general-purpose dataset, but would require the model to implement guardrails that would disable it from generating non-parodic imitations of individual authors’ oeuvres (?, ?, ?). [^12] As examples of guardrails, AI developers have implemented “refusal protocols” blocking outputs when prompts request content “in the style of” specific authors. Further, current reinforcement learning techniques can be easily modified to steer models away from stylistic imitation. Another solution, particularly where the lower quality of in-context prompting reduces the prospect of market dilution, might be to condition a ruling of fair use on the prominent disclosure of the output’s AI origin. This solution assumes that the public, informed that the output was not human-authored, will be less inclined to select the AI substitute; transparency should diminish the competition between human-authored and machine-generated offerings. The solution also assumes that in-context prompting will not in the future produce outputs that readers will prefer to human-authored text. Improvements in LLMs (and/or increases in the number of works copied into training data) may come to belie that assumption.
4 Methods
We recruited 28 candidates from top MFA programs (Iowa Writers’ Workshop, Helen Zell Writers’ Program at the University of Michigan, MFA Program in Creative Writing at New York University, Columbia University School of the Arts), paying each \$75 for writing a single excerpt. These MFA candidates and LLMs emulated the style/voice of 50 award-winning authors representing diverse cultural backgrounds and distinct literary voices (full list in Table 2 in SI). The writing prompt provided to MFA candidates contained (i) 20 sample excerpts spanning an author’s complete body of work (ii) textual descriptions of the author’s distinctive style and voice (iii) detailed content specifications about the original author written excerpt to be emulated. The selection of the 50 authors and all the writing prompts were developed in collaboration with five English Literature PhD students who analyzed each author’s literary voice and created the verbalized style descriptions (prompt in Figure 5 in SI).
[…]
These <<Human-AI>> pairs for both conditions were evaluated by 28 experts (the same MFA candidates) and 131 lay readers recruited from Prolific, [^13] https://www.prolific.com/ one of the leading crowdsourcing platforms for research participants. Experts never evaluated their own excerpts. Each pair was assessed by three experts and five lay readers, with majority voting determining final judgments. Inter-rater agreement was quantified using Fleiss’ kappa. In total, we obtained 2,400 pairwise evaluations (1,200 quality, 1,200 style) for AI Condition 1 and 1,440 evaluations (720 quality, 720 style) for AI Condition 2. For quality evaluation, we showed the <<Human-AI>> pair alone; for style evaluation, we included the original author written excerpt alongside the pair (see Figures 9 and 10 in SI). Both experts and lay readers also provided 2-3 sentence explanations grounded in textual evidence to justify their choices (?) (see Figures 1720 in SI).
To ensure annotation quality, we implemented attention checks including timestamp recording to prevent rushing. Additionally, we screened responses using Pangram, [^14] https://www.pangram.com/ a state-of-the-art AI detection tool, and excluded participants who used generative AI in their responses (?). Our study was approved by the University of Michigan IRB (HUM00264127) and was preregistered at OSF. [^15] https://osf.io/zt4ad Informed consent was obtained from all participants.
We tested hypotheses H1 (baseline LLMs vs. human writers) and H2 (fine-tuned GPT-4o vs. human writers) using logistic regression with CR2 cluster-robust standard errors clustered by reader. For H1, we compared human writing to the average performance across GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro in the in-context prompting condition. For H2, we directly compared fine-tuned GPT-4o to human writing. Contrasts were computed separately for expert and lay readers within each outcome (style, quality), with Holm correction applied across reader-group contrasts within each hypothesis-outcome combination. For H3 (AI detection and preference), we modeled the relationship between Pangram AI-detection scores and preference, testing whether fine-tuning attenuated detection-based penalties via setting $\times$ detection score interactions. All the model specifications are described in detail in SI.
Limitations
Our recruitment was mostly restricted to American creative writing programs and further study needs to be done across creative writing programs outside the US. While our pre-dedicated pool of 50 writers consisted of some writers who do not write in English, our experiments on style/voice emulation were done based on their English translation. Creative writing often depends on intrinsic motivation. While we offered MFA students a lucrative rate for writing the excerpts, it’s unclear if monetary incentives actually enhanced their creative output, since intrinsic motivation typically drives the best artistic work. Last but not least our experiments were conducted at a shorter excerpt level and conclusions cannot be drawn for long form text. In it’s current form AI is unable to generate long form text that’s thematically coherent unlike humans. While we foresee a situation where human can collaborate with a finetuned AI model to create competing long form works, experimental evidence is required to make any broader claims.
Code and Data availability
All analysis and figure-generation code alongside human data is available upon request. Core analyses can be reproduced by running the numbered R scripts in sequence. Analyses were conducted in R 4.3.1 with key packages including clubSandwich and emmeans. Exact package versions and run instructions are provided in SI, Section S10.
Ethics statement
All procedures involving human participants were approved by the University of Michigan Institutional Review Board (HUM00264127). Informed consent was obtained from all participants, who were compensated for their time. The study was preregistered at OSF (https://osf.io/zt4ad).
Acknowledgements
We thank Jared Brent Harbor (Columbia Law School, J.D. Class of 2027; M.F.A. in Theatre Management and Producing, Columbia University School of the Arts) for research and editorial assistance.
Author contributions
TC, PSD: Conceptualization; Methodology; Data Analysis; Writing (original draft); Writing (review and editing).
JCG: Writing (original draft); Writing (review and editing).
Competing interests
The authors declare no competing interests.
Materials & correspondence
Correspondence should be addressed to Tuhin Chakrabarty (tchakrabarty@cs.stonybrook.edu), Jane C. Ginsburg (ginsburg@law.columbia.edu), or Paramveer Dhillon (dhillonp@umich.edu).
Supplementary Information for
Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers
Tuhin Chakrabarty1, Jane C. Ginsburg2, Paramveer Dhillon3,4
1Department of Computer Science, Stony Brook University.
2Columbia Law School.
3School of Information Science, University of Michigan.
4MIT Initiative on the Digital Economy.
Corresponding authors: tchakrabarty@cs.stonybrook.edu, ginsburg@law.columbia.edu,
dhillonp@umich.edu
Materials and Methods
S1: Details about Writing Task
S1.1 Author List
Table 2 shows the list of 50 authors that were chosen by English Literature Ph.D. students. These chosen author list consists of canon-plus-global giants (Ernest Hemingway, Virginia Woolf, William Faulkner, Gabriel Garcia Márquez, Stephen King, Haruki Murakami, Kazuo Ishiguro), a strong set of contemporary prominent literary voices (Margaret Atwood, Ian McEwan, Jonathan Franzen, Colson Whitehead, George Saunders, Louise Erdrich, Octavia Butler, Salman Rushdie, Maya Angelou, Percival Everett), and a group of critically acclaimed / emerging authors (Ottessa Moshfegh, Tony Tulathimutte, Roxane Gay ). Additionally our author list is culturally diverse where several of the authors write primarily in a non English language (Han Kang, Yoko Ogawa, Annie Ernaux). Roughly two‑thirds (34) have secured at least one major international or national prize (e.g., Nobel, Booker, Pulitzer, National Book Award, MacArthur Fellowship, Women’s Prize, International Booker, Hugo/Nebula). 8 of the authors are Nobel Prize winners in Literature and 8 are Pulitzer Prize winners.
Table 1: List of Authors
# Author # Author # Author
1 Alice Munro 19 J.D. Salinger 37 Philip Roth
2 Annie Ernaux (✓) 20 Jhumpa Lahiri (✓) 38 Rachel Cusk
3 Annie Proulx (✓) 21 Joan Didion 39 Roxane Gay (✓)
4 Ben Lerner (✓) 22 Jonathan Franzen (✓) 40 Sally Rooney (✓)
5 Charles Bukowski 23 Junot Díaz (✓) 41 Salman Rushdie (✓)
6 Cheryl Strayed (✓) 24 Kazuo Ishiguro (✓) 42 Shirley Jackson
7 Chimamanda Ngozi Adichie (✓) 25 Louise Erdrich (✓) 43 Sigrid Nunez (✓)
8 Colson Whitehead (✓) 26 Lydia Davis (✓) 44 Stephen King
9 Cormac McCarthy 27 Margaret Atwood (✓) 45 Tony Tulathimutte (✓)
10 David Foster Wallace 28 Marilynne Robinson (✓) 46 V. S. Naipaul
11 Ernest Hemingway 29 Maya Angelou 47 Virginia Woolf
12 Flannery O’Connor 30 Milan Kundera 48 William Faulkner
13 Gabriel García Márquez 31 Min Jin Lee (✓) 49 Yoko Ogawa (✓)
14 George Saunders (✓) 32 Nora Ephron 50 Zadie Smith (✓)
15 Han Kang (✓) 33 Octavia Butler
16 Haruki Murakami (✓) 34 Orhan Pamuk (✓)
17 Hunter S. Thompson 35 Ottessa Moshfegh (✓)
18 Ian McEwan (✓) 36 Percival (✓)Everett
Table 2: Author list for our pool of 50 authors. (✓) denotes authors who were used in fine-tuning experiment
S1.2 Writing Prompt
Figure 5: In-Context Writing Prompt used in AI Condition 1.
Larger context windows in LLMs models allow for processing significantly more information at once, leading to improved accuracy, understanding, and complex reasoning capabilities. Taking advantage of this feature in GPT4-o, Claude-3.5-Sonnet and Gemini-1.5-Pro, we design long context prompt that first demonstrates 20 sample excerpts written by a given author, followed by their style/voice verbalized in text and finally the content of the excerpt to be emulated (See Figure 5). The same prompt was provided to both experts( MFA candidates) and LLMs.
S1.3 Finetuning details
Figure 6: The pipeline used to finetune ChatGPT on an authors entire oeuvre
For finetuning we bought digital ePub versions of these authors’ books and transformed them into plain text files.If the epub file was basically a wrapper around scanned page images, then we ignored that epub.The number of books written by each author vary a lot. For instance Tony Tulathimutte has written only two books Private Citizens and Rejection so we could only finetune GPT4-o on two of them. While for Haruki Murakami we could finetune on 22 books A Wild Sheep Chase, After Dark, After the Quake, Blind Willow, Sleeping Woman, Colorless Tsukuru Tazaki and his Years of Pilgrimage, Dance Dance Dance, First Person Singular, Hard-boiled Wonderland and the End of the World, Hear the Wind Sing, Kafka on the Shore, Killing Commendatore, Men Without Women, Norwegian Wood, Novelist as a Vocation, One and Two, Pinball, 1973, South of the Border, West of the Sun, Sputnik Sweetheart, The City and Its Uncertain Walls, The Elephant Vanishes, The Wind-Up Bird Chronicle, What I Talk About When I Talk About Running, Wind/Pinball: Two Novels
Finetuning on books isn’t a straightforward process. Each book is typically 50,000 to 80,000 words long with some exceptions. Using such long sequences wastes capacity because models still struggle with very long inputs and long‑context adaptation is non‑trivial; To preserve general capabilities in Supervised Finetuning it’s generally advised to have diverse, shorter samples. Breaking a book into context‑independent excerpts increases batch diversity and the number of distinct gradient signals per token budget. At the same time it also reduces overfitting to a single narrative flow while still injecting the salient stylistic/local patterns.
Our entire finetuning pipeline can be seen in Figure 6. We first convert the epub files to txt using https://github.com/kevinboone/epub2txt2. We then segment the entire book into context independent excerpts. At a first pass we naively split the entire book text by existing double-newlines and rejoin them to enforce excerpt size bounds(250-650 words). For the rare cases where the naive splitting would lead to excerpts longer than 650 words we used GPT4o to segment them again using the prompt Segment it into excerpts of minimum length 300-350 words such that each excerpt is grammatical from the start and doesn’t feel abruptly cut off. There should be zero deletion and break into excerpts at grammatically natural places. Maintain the original word count. Avoid breaking into too many small excerpts. Start directly. Don’t say Here’s or Here is ….
After obtaining context excerpts we extract content details from them using GPT4-o by prompting it Describe in detail what is happening in this excerpt. Mention the characters and whether the voice is in first or third person for majority of the excerpt.Maintain the order of sentences while describing.. Figure 7 shows a sample paragraph from Shame written by Annie Ernaux and the extracted content. Once we obtain the extracted content we finetune GPT4-o using their fine-tuning API with the following input-output pair Write a [[n]] word excerpt about the content below emulating the style and voice of [[authorname]]: [[excerpt]] as seen in Figure 6. This technique is commonly referred to as instruction back-translation (?)
Figure 7: Original author written excerpt and extracted content using an LLM
After finetuning is completed at inference time we can then simply generate an excerpt conditioned on a similar instruction that contains a novel content. We specifically excluded the original author written excerpt used for our experiments during fine-tuning so that the model does not get an unfair advantage. We also ensured that the generated outputs do not contain any memorized snippets from the original author written excerpt. In the rare occasion that a model regurgitated verbatim snippets/ngrams from original author written text we resampled it and manually verified that there is no verbatim overlap before using it for evaluation. While supervised finetuning in most cases ensures that the generated output contains all the information in the extracted content details, in the rare occasion that it does not we resample/regenerate it again. Last but not least sometimes supervised finetuning can lead to ungrammatical output or output with minor inconsistencies. To make sure these mistakes don’t impact human evaluation, we performed a post processing step using GPT4-o using the following prompt Fix grammar, tense, typo, spelling or punctuation error or any other awkward construction/ logical inconsistency
Figure 8: Total training token distribution for each author
For Margaret Atwood, Kazuo Ishiguro, Salman Rushdie and Haruki Murakami we fine-tuned GPT4-o for 1 epoch as they had multiple books/longer books (For instances . For the rest of the authors we fine-tuned for 3 epochs. We use default finetuning parameters set by OpenAI in addition to setting batch size as 1 and LR multiplier as 2. Table 16 shows that finetune models don’t regurgitate a lot of verbatim text from the training corpus (author’s entire oeuvre). We should also note that some of these words are also part of the Content in the instruction which would anyway appear in the text and should not be penalized. For all generation(prompting or fine-tuning default temperature=1.0 was used)
S2: Details about Evaluation Task
S2.1 Expert and Lay Recruitment Demographic
All experts(MFA candidates) recruited were currently residing in USA. This was a requirement for payment purposes as depending on how much an individual makes as a part of the study requires them to declare it as taxable income. Paying experts who do not have a ITIN or SSN would be challenging for logistical reasons. 65% of our expert did not identify as cis gender men. In terms of ethnicities our experts identified as South Asian, White, East Asian and Black. For lay readers recruited from Prolific we restricted ourselves to English speaking countries (only USA and UK). We also required participants to be born there, have a 100% acceptance rate and for everyone to be college educated.
S2.2 Evaluation Interface
Figure 9: Evaluation Interface for writing quality evaluation where readers (expert and lay) just see two excerpts without any disclaimer of its source and choose their preference supported by a reason
Figures 9 and 10 show the evaluation interface shown to both lay and expert readers. Readers read two excerpts for quality evaluation and three excerpts for stylistic fidelity evaluation given that the stylistic emulation is evaluated with respect to the original author written excerpt.
Figure 10: Evaluation Interface for stylistic fidelity evaluation where readers (expert and lay) just see three excerpts (original and two emulations) without any disclaimer of its source and choose their preference supported by a reason
Without observing why a choice was made (no rationale/side information), the latent factors driving that preference are often unclear. Such hidden user context, demographic/value variation, task ambiguity, or near‑tie similarity—remain inflates effective label noise. Providing rationales or auxiliary side information has been empirically shown to improve label reliability, data efficiency, and debiasing (?). This led us to ask for reasons that support the preference choice. For lay readers, reasons also act as proxy for understanding if the reader is doing the task sincerely and not choosing preference randomly.
S2.3 How much is faithful style/voice emulation important for good writing?
Figure 11: Pearson Correlation Coefficient between writing quality and style judgments
As can be seen in Figure11 writing quality has a strong correlation with a faithful style/voice fidelity for In-context Prompting condition compared to Finetuning. The high expert correlation for In-context Prompting suggests that models cannot emulate an author’s style properly just by prompting and that experts use stylistic cues as a major factor in quality judgments - likely detecting “AI-ness” (clichés, purple prose, too much exposition, lack of subtext, mixed metaphors) in the style that makes them rate quality lower too. With finetuning, experts can appreciate quality independent of stylistic fidelity, suggesting the as AI loses its tell-tale signs experts can evaluate each dimension on its own merits. For lay readers the correlation only drops slightly from 0.52 (In-context prompting) to 0.45 (fine-tuned), compared to experts’ dramatic drop from 0.64 to 0.39. This also mean that lay readers might focus on more surface-level features such as grammar, clarity or flow rather than subtle stylistic markers that distinguishes AI from human writing.
S2.4 Example of emulations
Figure 12,13, 14 and 15 shows emulations of the original excerpt written by Colson Whitehead, Ottessa Moshfegh, Jonathan Franzen and Han Kang. There are 3 distinct MFA emulations per author (MFA1, MFA2, MFA3). For the In-context Prompting set up we prompt Claude3.5 Sonnet (AI1), Gemini1.5 (AI2) and GPT-4o(AI3) to produce excerpts that can be pitted against MFA written excerpts. In Finetuning setup we finetune GPT4-o and compare the fine-tuned output AI4 against MFA1, MFA2, MFA3. The output from fine-tuned GPT4-o (AI4) was generated using default hyperparameters.
Figure 12: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 13: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 14: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 15: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 16: Total overlap percentage calculated by number of words in generated output that come from authors overall corpus
S2.5 Dissecting Preference Evaluations from Experts
Figure 17 and 18 shows the writing quality preference evaluation results from expert readers. Expert preferences are often grounded in solid reasoning and they often agree on similar snippets to support their reasoning. For instance look at uncooked spaghetti as an overwrought/awkward metaphor highlighted by both Expert Reader 1 and 3 while tarot card predicting his fate highlighted by Expert Reader 2 and 3. Grounding preferences in reasoning helps us uncover their mental models. It is also worth noticing that the preference often shifts towards AI once it is fine-tuned on authors entire oeuvre. Fine-tuning helps get rid of the officious disembodied robovoice that is characteristic of ChatGPT. Here experts praise the model for idiosyncratic narrative voice. By default due to post training guardrails GPT4-o generates a rather safe and somewhat polite text in the in-context prompting set up. But what is particularly impressive here is when fine-tuned the vocabulary of the model generated text veers towards ribald and somewhat profane text that is characteristic of Tulathimutte’s voice (particularly Rejection from where the text is drawn). In Figure 19 we see that experts prefer the MFA emulation as more faithful to Junot Diaz’s voice. In fact the excerpt written by Gemini-1.5-Pro is oddly poetic and rife with purple prose and cliches which is nothing like Junot Diaz’s voice that is often marked by a dazzling hash of Spanish, English, slang, literary flourishes, and pure virginal dorkiness as noted by all three expert readers. However when fine-tuned on authors’ entire oeuvre (Figure 20) we see that the model learns these references and uses them in a rather wry and humorous manner (papi chulo, fifty pendejas, little pito off, las plagas) that is appreciated by expert readers leading them to prefer AI over MFAs.
Figure 17: MFA vs AI emulation of Tony Tulathimutte, readerd by 3 expert readers all of whom agree that MFA written excerpt is superior in terms of writing quality.
Figure 18: MFA vs AI emulation of Tony Tulathimutte, judged by 3 expert readers all of whom agree that Fine-tuned AI written excerpt is superior in terms of writing quality
Figure 19: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers all of whom agree that MFA written excerpt is superior in terms of stylistic fidelity.
Figure 20: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers all of whom agree that Fine-tuned AI written excerpt is superior in terms of stylistic fidelity
S3: Details about AI detection and Stylometry
S3.1 AI detection thresholds
Figure 21: Pangram AI likelihood scores vs GPTZero AI Likelihood scores for MFA, In-context and Fine-tuned
Based on AI likelihood scores from Pangram and GPTZero across 330 (150 MFA, 150 In-context AI, 30 Fine-tuned AI) evaluations we see that Pangram is bimodal. This means that AI likelihood scores are 0 most of the times for human written text. While there are few spikes, a conservative threshold of 0.9 drastically reduces any chance of a False Positive. Similarly, AI likelihood scores are 1.0 most of the times for AI written text with few exceptions. GPTZero is less accurate compared to Pangram, however a threshold of 0.9 holds as a good threshold for it too. GPTZero considers fine-tuned AI generation to be more human like compared to Pangram.
S3.2 Calculating Cliché Density
Figure 22: Prompt to identify Clichés
Language models are good at identifying specific patterns. Taking advantage of this we prompt Claude 4.1 Opus to generate list of clichés given an excerpt. Figure 22 shows the prompt used to identify Clichés. However, sometimes language models can overgenerate and penalize expressions that are not clichés. To address this, two authors of the papers separately annotated which of the expressions are actually clichés from the list. We take intersection of their individual list as the final list of clichés per paragraph. To calculate Cliché Density we then used following formula
[…]
S4: Statistical Modeling and Analysis
S4.1: Data Structure and Notation
The analysis employs trial-level data in long format where each row represents a reader’s preference decision for a single excerpt within a pairwise comparison. Each pairwise judgment contributes two rows (one per excerpt with $i\in\{1,2\}$ sharing the same $j,k$). Let $Y_{ijk}\in\{0,1\}$ denote the preference outcome where $Y_{ijk}=1$ if excerpt $i$ is preferred by reader $j$ in comparison $k$, and 0 otherwise. The dataset comprises:
  • In-context prompting: 2,400 judgments (1,200 per outcome); 4,800 long-format rows.
  • Fine-tuning: 1,440 judgments (720 per outcome); 2,880 long-format rows.
Key variables include writer type $W_{i}\in\{\text{Human},\text{GPT-4o},\text{Claude},\text{Gemini},\text{GPT-4o-FT}\}$, reader type $J_{j}\in\{\text{Expert},\text{Lay}\}$, and for H3, the Pangram AI-detection score $s_{i}\in[0,1]$. Throughout, Human and Expert serve as reference categories unless otherwise specified. Standard errors use CR2 with readers as the clustering unit; rows within a judgment pair are not independent, but in this design clustering by reader (the source of repeated measures) is the dominant dependence.
S4.2: Primary Models for H1 and H2
We test our preregistered hypotheses using fixed-effects logistic regression with CR2 cluster-robust standard errors. The CR2 estimator provides improved finite-sample performance compared to standard cluster-robust estimators, particularly important given our relatively small number of readers (28 experts, 131 lay readers).
S4.2.1: H1: Baseline LLMs vs Human (In-context Setting)
For H1, we analyze only the in-context prompting trials containing Human and baseline LLMs (GPT-4o, Claude, Gemini), excluding fine-tuned excerpts. We fit the logistic regression:
$\text{logit}\,P(Y_{ijk}=1)=\alpha+\sum_{m\in\mathcal{M}}\beta_{m}\mathbf{1}[W_{i}=m]+\gamma\mathbf{1}[J_{j}=\text{Lay}]+\sum_{m\in\mathcal{M}}\phi_{m}\mathbf{1}[W_{i}=m]\cdot\mathbf{1}[J_{j}=\text{Lay}]$ (2)
where $\mathcal{M}=\{\text{GPT-4o},\text{Claude},\text{Gemini}\}$ and Human serves as the reference category. The interaction terms $\phi_{m}$ capture differential preferences between expert and lay readers. The preregistered H1 contrast tests whether humans outperform the average of baseline LLMs:
$\Delta^{\text{H1}}_{g}=\frac{1}{3}\sum_{m\in\mathcal{M}}\eta(m,g)-\eta(\text{Human},g)$ (3)
where $\eta(w,g)$ denotes the linear predictor (fitted log-odds) for writer $w$ and reader group $g\in\{\text{Expert},\text{Lay}\}$. The odds ratio $\text{OR}^{\text{H1}}_{g}=\exp(\Delta^{\text{H1}}_{g})$ quantifies the relative preference, with values <1<1 indicating human preference and values >1>1 indicating AI preference.
S4.2.2: H2: Fine-tuned GPT-4o vs Human
For H2, we analyze only the fine-tuning trials containing Human and GPT-4o-FT excerpts. We fit:
$\text{logit}\,P(Y_{ijk}=1)=\alpha+\beta_{\text{FT}}\mathbf{1}[W_{i}=\text{GPT-4o-FT}]+\gamma\mathbf{1}[J_{j}=\text{Lay}]+\phi_{\text{FT}}\mathbf{1}[W_{i}=\text{GPT-4o-FT}]\cdot\mathbf{1}[J_{j}=\text{Lay}]$ (4)
The H2 contrast directly compares fine-tuned GPT-4o to human writers:
$\Delta^{\text{H2}}_{g}=\eta(\text{GPT-4o-FT},g)-\eta(\text{Human},g)$ (5)
yielding $\text{OR}^{\text{H2}}_{g}=\exp(\Delta^{\text{H2}}_{g})$. This tests whether fine-tuning achieves parity ($\text{OR}\approx 1$) or superiority ($\text{OR}>1$) compared to human experts.
All contrasts and confidence intervals use Wald (normal) inference with the CR2 covariance matrix. Specifically, 95% CIs are computed as estimate $\pm 1.96\cdot$ SE on the log-odds scale, then exponentiated. Holm correction is applied across the two reader-group contrasts (Expert, Lay) within each outcome (style, quality) and hypothesis (H1, H2).
S4.3: H3: AI Detection and Preference
To test whether AI detectability influences preferences and whether fine-tuning removes this relationship, we model:
$\text{logit}\,P(Y_{ijk}=1)=$ $\alpha+\beta_{1}s_{i}+\beta_{2}\mathbf{1}[\text{Setting}_{i}=\text{FT}]+\beta_{3}\mathbf{1}[J_{j}=\text{Lay}]$
$+\beta_{12}s_{i}\cdot\mathbf{1}[\text{Setting}_{i}=\text{FT}]+\beta_{13}s_{i}\cdot\mathbf{1}[J_{j}=\text{Lay}]$
$+\beta_{23}\mathbf{1}[\text{Setting}_{i}=\text{FT}]\cdot\mathbf{1}[J_{j}=\text{Lay}]$
$+\beta_{123}s_{i}\cdot\mathbf{1}[\text{Setting}_{i}=\text{FT}]\cdot\mathbf{1}[J_{j}=\text{Lay}]$ (6)
where $s_{i}$ is the Pangram score (continuous, 0-1) and Setting indicates in-context vs fine-tuned. The coefficient $\beta_{1}$ captures the detection penalty in the baseline condition—how much AI detectability reduces preference. The interaction term $\beta_{12}$ tests whether fine-tuning attenuates this relationship, with a positive value indicating that fine-tuning removes the penalty associated with AI detection.
S4.4: Exploratory Analyses
S4.4.1: Stylometric Mediation
The mediation analysis examines which textual features explain the link between AI detection and human preferences. We follow a two-stage approach. First, we regress Pangram scores on stylometric features using elastic net regularization to handle multicollinearity:
$s_{i}=\alpha+\sum_{k}\lambda_{k}X_{ik}+\epsilon_{i}$ (7)
where $X_{ik}$ includes cliché density, sentence-length variance, and parts-of-speech proportions. Features surviving regularization are then included alongside Pangram scores in the preference model to decompose the total effect into direct and mediated components. The proportion mediated is calculated using the product-of-coefficients method.
S4.4.2: Author-Level Heterogeneity
To assess variability across the 30 fine-tuned authors, we compute empirical preference rates using Jeffreys prior estimates to stabilize small-sample authors:
$\hat{p}_{a}=\frac{k_{a}+0.5}{n_{a}+1}$ (8)
with 95% intervals from Beta($\frac{1}{2}$, $\frac{1}{2}$) priors. The relationship between preference rates and fine-tuning corpus size (in millions of tokens) is assessed via OLS regression with heteroskedasticity-robust standard errors.
S4.5: Mapping to Figures
The model outputs map directly to figures in the main text:
  • Figures 2A-B: Exponentiated contrasts $\text{OR}^{\text{H1}}_{g}$ and $\text{OR}^{\text{H2}}_{g}$ with 95% CIs computed on log-odds scale then transformed
  • Figures 2C-D: Predicted probabilities $p(w,g)=\text{logit}^{-1}(\eta(w,g))$ with delta-method confidence intervals
  • Figure 2F: Model-implied probabilities from H3 evaluated at $s\in\{0.1,0.3,0.5,0.7,0.9\}$
  • Figure 3: Author-level preference rates $\hat{p}_{a}$ plotted against corpus size with OLS regression lines
  • Figure 4A: Stylometric mediation analysis showing proportion of Pangram effect explained by textual features
  • Figure 4B: Fine-tuning premium (difference in AI preference between fine-tuned and in-context) versus corpus size
  • Figure 4C: Cost comparison for producing 100,000 words of text
Inter-rater agreement quantifies consistency beyond chance using Fleiss’ kappa, appropriate here as each pair was evaluated by multiple readers (3 experts and 5 lay readers):
$\kappa=\frac{\bar{P}-\bar{P}_{e}}{1-\bar{P}_{e}}$ (9)
where $\bar{P}$ represents mean observed agreement and $\bar{P}_{e}$ expected agreement by chance. Expert readers achieved moderate agreement ($\kappa=0.41-0.67$) while lay readers showed minimal agreement ($\kappa=0.07-0.22$), reflecting the subjective nature of literary evaluation.
Full code and exact package versions used to generate Figures 2-4 are provided in the Code & Data Availability section (S10) and the OSF repository.
S5.1: Cell Counts by Experimental Conditions
Unless noted otherwise, counts below refer to reader-level long-format observations entering the logistic models (two rows per judgment—one row per alternative in the pair). In the in-context prompting setting, 150 human–AI pairs were evaluated (evenly split across three baseline models: 50 Human vs GPT-4o, 50 Human vs Claude 3.5 Sonnet, 50 Human vs Gemini 1.5 Pro), yielding 1,200 judgments per outcome and 2,400 long-format rows. In the fine-tuned setting, 90 human–AI pairs yielded 720 judgments per outcome and 1,440 long-format rows. Human rows are three times more frequent than any single AI baseline in the in-context prompting setting because every pair contains a Human excerpt, while the 150 pairs are split evenly across the three baseline models.
Table 3: Observation counts for stylistic fidelity in the in-context prompting setting. Each row is a long-format observation entering the H1 analysis.
Outcome Setting Writer type Reader type $n$
style In_Context Human Expert 450
style In_Context Human Lay 750
style In_Context GPT4o_baseline Expert 150
style In_Context GPT4o_baseline Lay 250
style In_Context Claude_baseline Expert 150
style In_Context Claude_baseline Lay 250
style In_Context Gemini_baseline Expert 150
style In_Context Gemini_baseline Lay 250
Table 4: Observation counts for stylistic fidelity in the fine-tuned setting. Balanced counts (Human = GPT4o_fine-tuned) reflect the one-to-one comparison design for H2.
Outcome Setting Writer type Reader type $n$
style Fine_tuned Human Expert 270
style Fine_tuned Human Lay 450
style Fine_tuned GPT4o_fine-tuned Expert 270
style Fine_tuned GPT4o_fine-tuned Lay 450
Table 5: Observation counts for writing quality in the in-context prompting setting (same Human:AI baseline ratio as stylistic fidelity).
Outcome Setting Writer type Reader type $n$
quality In_Context Human Expert 450
quality In_Context Human Lay 750
quality In_Context GPT4o_baseline Expert 150
quality In_Context GPT4o_baseline Lay 250
quality In_Context Claude_baseline Expert 150
quality In_Context Claude_baseline Lay 250
quality In_Context Gemini_baseline Expert 150
quality In_Context Gemini_baseline Lay 250
Table 6: Observation counts for writing quality in the fine-tuned setting.
Outcome Setting Writer type Reader type $n$
quality Fine_tuned Human Expert 270
quality Fine_tuned Human Lay 450
quality Fine_tuned GPT4o_fine-tuned Expert 270
quality Fine_tuned GPT4o_fine-tuned Lay 450
S5.2: Writer Category Mapping
Table 7 provides the mapping from raw data labels to the analysis categories used in Section S4. This mapping ensures reproducibility and clarifies the distinction between baseline (in-context) and fine-tuned AI conditions.
Table 7: Mapping of writer categories from raw data fields to analysis labels.
Writer category Description Mapping rule (from raw fields)
Human Human-written (MFA authors) excerpt_type = Human
GPT4o_baseline GPT-4o in-context prompting AI excerpt; excerpt_model $\in$ {GPT4o}; setting = In_Context
Claude_baseline Claude 3.5 Sonnet (in-context) AI excerpt; excerpt_model = Claude3.5Sonnet; setting = In_Context
Gemini_baseline Gemini 1.5 Pro (in-context) AI excerpt; excerpt_model = Gemini_1.5_Pro; setting = In_Context
GPT4o_fine-tuned GPT-4o fine-tuned AI excerpt; excerpt_model = GPT4o_Fine-tuned
S6: Main GLM Results and Contrasts (Figure 2)
This section presents the numerical results from the logistic regression models specified in Section S4. All models use CR2 cluster-robust standard errors with readers as the clustering unit. The reference category throughout is Human × Expert. These tables provide the complete statistical foundation for Figure 2 in the main text.
S6.1: Model Coefficients
Table 8 reports the full coefficient tables for all four primary models. The dramatic sign reversal between in-context prompting and fine-tuned settings is immediately apparent: negative coefficients for AI writers in in-context prompting (indicating human preference) become positive in fine-tuning (indicating AI preference). The writer type $\times$ reader type interactions in the in-context prompting models reveal that lay readers are substantially more favorable to AI-generated text than experts, a difference that disappears after fine-tuning.
Table 8: GLM coefficients with CR2 robust standard errors for all primary models. Negative coefficients indicate preference for the reference category (Human), while positive coefficients indicate preference for AI. The sign reversal between in-context prompting and fine-tuned models represents the core finding.
Term Est. (log-odds) SE $z$ $p$
Style — In-context Prompting (n = 2,400)
(Intercept) 0.901 0.148 6.098 1.07e-09
writer_typeGPT4o_baseline -1.655 0.402 -4.120 3.78e-05
writer_typeClaude_baseline -1.390 0.276 -5.039 4.67e-07
writer_typeGemini_baseline -2.510 0.426 -5.894 3.77e-09
reader_typeLay -0.826 0.169 -4.877 1.08e-06
writer_typeGPT4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
writer_typeClaude_baseline:reader_typeLay 1.428 0.336 4.249 2.15e-05
writer_typeGemini_baseline:reader_typeLay 2.211 0.475 4.650 3.32e-06
Style — Fine-tuned (n = 1,440)
(Intercept) -1.050 0.141 -7.435 1.04e-13
writer_typeGPT4o_fine-tuned 2.100 0.282 7.435 1.04e-13
reader_typeLay -0.008 0.186 -0.042 0.967
writer_typeGPT4o_fine-tuned:reader_typeLay 0.015 0.372 0.042 0.967
Quality — In-context Prompting (n = 2,400)
(Intercept) 0.978 0.174 5.634 1.76e-08
writer_typeGPT4o_baseline -2.406 0.473 -5.083 3.71e-07
writer_typeClaude_baseline -1.246 0.368 -3.390 6.997e-04
writer_typeGemini_baseline -2.406 0.421 -5.711 1.13e-08
reader_typeLay -1.197 0.196 -6.124 9.12e-10
writer_typeGPT4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
writer_typeClaude_baseline:reader_typeLay 1.723 0.420 4.105 4.04e-05
writer_typeGemini_baseline:reader_typeLay 2.594 0.473 5.480 4.25e-08
Quality — Fine-tuned (n = 1,440)
(Intercept) -0.314 0.122 -2.571 0.0102
writer_typeGPT4o_fine-tuned 0.627 0.244 2.571 0.0102
reader_typeLay -0.129 0.157 -0.821 0.412
writer_typeGPT4o_fine-tuned:reader_typeLay 0.258 0.314 0.821 0.412
S6.2: Primary Hypothesis Tests
Table 9 presents the preregistered contrasts testing H1 and H2, corresponding to panels A-B in Figure 2. The odds ratios reveal a striking reversal: expert readers’ 6-8 fold preference for human writing in in-context prompting conditions (OR = 0.16 for style, 0.13 for quality) transforms into an 8-fold preference for AI in fine-tuning (OR = 8.16 for style). Lay readers show less dramatic but directionally consistent shifts.
Table 9: Primary contrasts testing H1 (baseline LLMs vs Human) and H2 (fine-tuned vs Human). Holm correction applied within each outcome and hypothesis across reader types. The OR column shows odds ratios with values <1<1 favoring humans and >1>1 favoring AI.
Outcome Hyp. Reader Est. SE $p$ $p_{\mathrm{Holm}}$ OR 95% CI
style H1 Expert -1.852 0.308 3.16e-09 1.26e-08 0.157 [0.085, 0.290]
style H1 Lay -0.150 0.166 0.365 0.365 0.861 [0.621, 1.193]
style H2 Expert 2.101 0.276 2.19e-14 2.19e-14 8.163 [4.693, 14.198]
style H2 Lay 2.116 0.241 1.76e-18 1.76e-18 8.290 [5.155, 13.333]
quality H1 Expert -2.021 0.405 1.16e-07 2.33e-07 0.133 [0.063, 0.280]
quality H1 Lay 0.441 0.180 0.014 0.019 1.554 [1.092, 2.212]
quality H2 Expert 0.626 0.248 0.012 0.016 1.873 [1.161, 3.021]
quality H2 Lay 0.886 0.197 7.08e-06 1.42e-05 2.424 [1.644, 3.574]
S6.3: Model-Predicted Probabilities
Table 10 reports the predicted probabilities displayed in Figure 2 panels C-D. These probabilities, derived from the inverse-logit transformation of linear predictors, show the convergence of expert and lay preferences after fine-tuning. While experts strongly prefer humans over all baseline models in in-context prompting settings (probabilities 0.17-0.43), fine-tuning elevates AI preference to approximately 0.74 for both reader groups.
Table 10: Predicted probabilities of selecting AI excerpts by model and reader type, corresponding to Figure 2C-D. Values above 0.5 indicate AI preference. Note the convergence of expert and lay preferences in fine-tuned models.
Outcome Setting Model Reader $\hat{p}$ 95% CI
style In_Context GPT-4o (In-Context) Expert 0.320 [0.214, 0.449]
style In_Context GPT-4o (In-Context) Lay 0.472 [0.407, 0.538]
style In_Context Claude (In-Context) Expert 0.380 [0.303, 0.464]
style In_Context Claude (In-Context) Lay 0.528 [0.464, 0.591]
style In_Context Gemini (In-Context) Expert 0.167 [0.097, 0.271]
style In_Context Gemini (In-Context) Lay 0.444 [0.374, 0.516]
style Fine_tuned GPT-4o (Fine-tuned) Expert 0.741 [0.684, 0.790]
style Fine_tuned GPT-4o (Fine-tuned) Lay 0.742 [0.694, 0.785]
quality In_Context GPT-4o (In-Context) Expert 0.193 [0.112, 0.313]
quality In_Context GPT-4o (In-Context) Lay 0.608 [0.541, 0.671]
quality In_Context Claude (In-Context) Expert 0.433 [0.329, 0.544]
quality In_Context Claude (In-Context) Lay 0.564 [0.499, 0.627]
quality In_Context Gemini (In-Context) Expert 0.193 [0.124, 0.289]
quality In_Context Gemini (In-Context) Lay 0.492 [0.422, 0.563]
quality Fine_tuned GPT-4o (Fine-tuned) Expert 0.578 [0.519, 0.635]
quality Fine_tuned GPT-4o (Fine-tuned) Lay 0.609 [0.562, 0.654]
S6.4: Interaction Diagnostics
Tables 11 and 12 examine the writer type × reader type interactions in detail. The significant interactions in in-context prompting models (all $p<0.001$) quantify lay readers’ greater tolerance for AI-generated text. The absence of significant interactions in fine-tuned models ($p>0.4$) indicates that fine-tuning produces text that both expert and lay readers find equally compelling.
Table 11: Individual interaction terms showing differential preferences between expert and lay readers. Large positive coefficients in in-context prompting models indicate lay readers are more favorable to AI than experts.
Outcome Setting Term Est. SE $z$ $p$
style In_Context GPT4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
style In_Context Claude_baseline:reader_typeLay 1.428 0.336 4.249 2.15e-05
style In_Context Gemini_baseline:reader_typeLay 2.211 0.475 4.650 3.32e-06
style Fine_tuned GPT4o_fine-tuned:reader_typeLay 0.015 0.372 0.042 0.967
quality In_Context GPT4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
quality In_Context Claude_baseline:reader_typeLay 1.723 0.420 4.105 4.04e-05
quality In_Context Gemini_baseline:reader_typeLay 2.594 0.473 5.480 4.25e-08
quality Fine_tuned GPT4o_fine-tuned:reader_typeLay 0.258 0.314 0.821 0.412
Table 12: Joint Wald tests for writer type × reader type interactions. Significant interactions in in-context prompting models become non-significant after fine-tuning, indicating convergence of expert and lay preferences.
Outcome Setting Term df $\chi^{2}$ $p$
style In_Context writer_type:reader_type 3 24.923 1.60e-05
style Fine_tuned writer_type:reader_type 1 0.0017 0.967
quality In_Context writer_type:reader_type 3 37.584 3.46e-08
quality Fine_tuned writer_type:reader_type 1 0.674 0.412
The results in this section provide compelling statistical evidence for the paper’s central finding: fine-tuning on author-specific corpora fundamentally transforms AI-generated text from clearly inferior (as judged by experts) to preferred over human writing. The convergence of expert and lay preferences after fine-tuning suggests that the improvements are not merely superficial but represent genuine advances in literary quality and stylistic fidelity.
S7: Author-Level Heterogeneity (Figure 3)
This section examines variation in AI preference rates across the 30 fine-tuned authors. Despite substantial differences in training corpus sizes (0.89M to 10.9M tokens) and fine-tuning costs (\$22 to \$273), we find no detectable relationship between data quantity and model performance within this token range, suggesting that even authors with limited published works can be effectively emulated.
S7.1: Per-Author Success Rates
Table LABEL:tab:s7.1 presents author-level AI preference rates using Jeffreys prior estimates to stabilize small-sample authors. The results reveal striking heterogeneity. For style, AI preference rates range from 18% (Tony Tulathimutte) to 98% (Roxane Gay), with 27 of 30 authors showing AI superiority (rate >> 0.5). For quality, the range spans 30% (Ian McEwan) to 86% (Cheryl Strayed), with 23 of 30 authors favoring AI.
Notably, some authors show divergent patterns across outcomes. Tony Tulathimutte represents an extreme case: readers found his style uniquely difficult to emulate (18% AI preference) yet rated AI quality as acceptable (54%). This suggests certain idiosyncratic voices resist algorithmic mimicry even when technical writing competence is achieved.
Table 13: Per-author AI preference rates (Jeffreys-smoothed and raw) ranked by performance. Values above 0.5 indicate AI preference. Note the wide variation across authors and the divergence between style and quality rankings for some authors.
Outcome Author Rank AI Win Rate (Jeffreys) Human Win Rate (Jeffreys) AI Win Rate (Raw)
quality Cheryl Strayed 1.0 0.86 0.14 0.875
quality Marilynne Robinson 2.0 0.78 0.22 0.792
quality Colson Whitehead 3.0 0.74 0.26 0.750
quality Han Kang 4.0 0.74 0.26 0.750
quality Haruki Murakami 5.0 0.74 0.26 0.750
quality Junot Diaz 6.0 0.74 0.26 0.750
quality Rachel Cusk 7.0 0.74 0.26 0.750
quality Salman Rushdie 8.0 0.74 0.26 0.750
quality Sigrid Nunez 9.0 0.74 0.26 0.750
quality Orhan Pamuk 10.0 0.70 0.30 0.708
quality Lydia Davis 11.0 0.66 0.34 0.667
quality Percival Everett 12.0 0.66 0.34 0.667
quality Jonathan Franzen 13.0 0.62 0.38 0.625
quality Louise Erdrich 14.0 0.62 0.38 0.625
quality Annie Proulx 15.0 0.58 0.42 0.583
quality George Saunders 16.0 0.58 0.42 0.583
quality Zadie Smith 17.0 0.58 0.42 0.583
quality Chimamanda Ngozi Adichie 18.0 0.54 0.46 0.542
quality Jhumpa Lahiri 19.0 0.54 0.46 0.542
quality Margaret Atwood 20.0 0.54 0.46 0.542
quality Roxane Gay 21.0 0.54 0.46 0.542
quality Sally Rooney 22.0 0.54 0.46 0.542
quality Tony Tulathimutte 23.0 0.54 0.46 0.542
quality Min Jin Lee 24.0 0.46 0.54 0.458
quality Annie Ernaux 25.0 0.42 0.58 0.417
quality Ben Lerner 26.0 0.42 0.58 0.417
quality Kazuo Ishiguro 27.0 0.42 0.58 0.417
quality Ottessa Moshfegh 28.0 0.38 0.62 0.375
quality Yoko Ogawa 29.0 0.34 0.66 0.333
quality Ian McEwan 30.0 0.30 0.70 0.292
style Roxane Gay 1.0 0.98 0.02 1.000
style Chimamanda Ngozi Adichie 2.0 0.94 0.06 0.958
style Han Kang 3.0 0.94 0.06 0.958
style Margaret Atwood 4.0 0.94 0.06 0.958
style Ben Lerner 5.0 0.90 0.10 0.917
style Junot Diaz 6.0 0.90 0.10 0.917
style Marilynne Robinson 7.0 0.90 0.10 0.917
style Kazuo Ishiguro 8.0 0.86 0.14 0.875
style Lydia Davis 9.0 0.86 0.14 0.875
style Orhan Pamuk 10.0 0.82 0.18 0.833
style Cheryl Strayed 11.0 0.78 0.22 0.792
style Min Jin Lee 12.0 0.78 0.22 0.792
style Sigrid Nunez 13.0 0.78 0.22 0.792
style Colson Whitehead 14.0 0.74 0.26 0.750
style Haruki Murakami 15.0 0.74 0.26 0.750
style Ian McEwan 16.0 0.74 0.26 0.750
style Rachel Cusk 17.0 0.74 0.26 0.750
style Sally Rooney 18.0 0.74 0.26 0.750
style Percival Everett 19.0 0.70 0.30 0.708
style Jhumpa Lahiri 20.0 0.66 0.34 0.667
style Ottessa Moshfegh 21.0 0.66 0.34 0.667
style Zadie Smith 22.0 0.66 0.34 0.667
style Annie Proulx 23.0 0.62 0.38 0.625
style George Saunders 24.0 0.62 0.38 0.625
style Jonathan Franzen 25.0 0.62 0.38 0.625
style Salman Rushdie 26.0 0.62 0.38 0.625
style Yoko Ogawa 27.0 0.62 0.38 0.625
style Louise Erdrich 28.0 0.50 0.50 0.500
style Annie Ernaux 29.0 0.42 0.58 0.417
style Tony Tulathimutte 30.0 0.18 0.82 0.167
Note. Jeffreys prior adds 0.5 to successes and 0.5 to failures; differences from raw are small but avoid 0/1 edge cases.
S7.2: Relationship with Corpus Size
Table 14 presents OLS regression results examining the relationship between fine-tuning corpus size (ranging from 0.89M tokens for Tulathimutte to 10.9M for Pamuk) and AI preference rates. The near-zero slopes with confidence intervals spanning zero and minimal $R^{2}$ values indicate no detectable relationship within this token range between training data quantity and model performance.
Table 14: OLS regression of author-level AI win rate on fine-tuning corpus size (millions of tokens). Slopes are near zero with CIs spanning zero and very low $R^{2}$, indicating no detectable relationship between corpus size and AI preference rate.
Outcome Slope Slope CI (Low) Slope CI (High) Intercept Intercept CI (Low) Intercept CI (High) $R^{2}$ N Authors
quality 0.0070 -0.0174 0.0314 0.570 0.458 0.681 0.0122 30
style 0.0032 -0.0262 0.0326 0.729 0.595 0.863 0.0017 30
Non-parametric check. Spearman $\rho=0.138$ ($p=0.467$) for quality and $\rho=-0.140$ ($p=0.462$) for style.
This finding has economic implications: authors with limited published works (costing \$22–\$50 to fine-tune) can be emulated as effectively as prolific authors (costing \$200+). The absence of any detectable corpus-size effect ($R^{2}\approx 0$) suggests that capturing an author’s voice depends more on stylistic consistency than sheer data volume.
S8: Detection Mechanisms and Economic Analysis (Figure 4)
This section examines the mechanisms linking AI detectability to human preferences (H3) and quantifies the economic implications of fine-tuning. We show that fine-tuning reverses and substantially attenuates the negative association between AI detection and preference observed in in-context prompting, while achieving performance that readers prefer at $\approx$99.7% lower cost than human writers.
S8.1: AI Detection and Preference Relationship
Table LABEL:tab:s8.1 presents the full model examining how AI detectability (Pangram score) influences preferences. The key finding is the significant interaction between pangram_score and setting (quality: $\hat{\beta}\approx 2.90$, $p<0.001$), indicating that fine-tuning reverses and substantially attenuates the negative association between detectability and preference observed in in-context prompting.
Table 15: Pangram GLM coefficients with CR2 robust standard errors (clustered by reader). Negative coefficients for pangram_score in in-context prompting indicate detection reduces preference; positive interactions with setting show that fine-tuning reverses and attenuates this penalty.
Outcome NN NN Readers Term Estimate SE $z$ $p$ OR OR (Low) OR (High)
quality 3840 159 (Intercept) 0.990 0.169 5.87 0.000 2.69 1.93 3.74
quality 3840 159 pangram_score -2.010 0.333 -6.03 0.000 0.13 0.07 0.26
quality 3840 159 setting (Fine-tuned) -1.020 0.175 -5.86 0.000 0.36 0.26 0.51
quality 3840 159 reader_type (Lay) -1.220 0.190 -6.42 0.000 0.30 0.20 0.43
quality 3840 159 pangram_score $\times$ setting 2.900 0.876 3.32 0.001 18.20 3.28 102.00
quality 3840 159 pangram_score $\times$ reader_type 2.480 0.377 6.56 0.000 11.90 5.67 24.90
quality 3840 159 setting $\times$ reader_type 1.250 0.198 6.33 0.000 3.50 2.37 5.15
quality 3840 159 pangram_score $\times$ setting $\times$ reader_type -3.290 1.030 -3.21 0.001 0.04 0.00 0.28
style 3840 159 (Intercept) 0.909 0.146 6.23 0.000 2.48 1.86 3.31
style 3840 159 pangram_score -1.850 0.292 -6.32 0.000 0.16 0.09 0.28
style 3840 159 setting (Fine-tuned) -0.938 0.158 -5.95 0.000 0.39 0.29 0.53
style 3840 159 reader_type (Lay) -0.832 0.167 -4.97 0.000 0.44 0.31 0.60
style 3840 159 pangram_score $\times$ setting 2.560 0.810 3.15 0.002 12.90 2.63 63.00
style 3840 159 pangram_score $\times$ reader_type 1.690 0.336 5.03 0.000 5.41 2.80 10.50
style 3840 159 setting $\times$ reader_type 0.840 0.178 4.73 0.000 2.32 1.64 3.28
style 3840 159 pangram_score $\times$ setting $\times$ reader_type -1.900 0.922 -2.06 0.039 0.15 0.02 0.91
Notes. CR2 robust standard errors clustered by reader. Readers: 28 Expert and 131 Lay (total $N_{\text{readers}}=159$).
S8.2: Predicted Probabilities Across Detection Levels
Tables 16 and 17 show model-predicted probabilities at different AI detection levels, corresponding to Figure 2F. For expert readers evaluating quality, high detection scores (0.9) reduce AI preference to $\sim$31% in in-context prompting but have a point estimate of $\sim$68% after fine-tuning (95% CI 0.370.370.890.89), demonstrating that fine-tuning makes outputs robust to detection-based skepticism.
Table 16: Predicted AI preference probabilities for style at different Pangram detection scores. Fine-tuning reverses and attenuates the relationship with detection score (with wider uncertainty at high detection).
Bin Setting Reader Type Prob Prob (Low) Prob (High)
0.1 In-context Prompting Expert 0.674 0.621 0.722
0.5 In-context Prompting Expert 0.497 0.494 0.500
0.9 In-context Prompting Expert 0.320 0.273 0.372
0.1 Fine-tuned Expert 0.511 0.491 0.530
0.5 Fine-tuned Expert 0.581 0.435 0.714
0.9 Fine-tuned Expert 0.648 0.380 0.847
0.1 In-context Prompting Lay 0.515 0.483 0.547
0.5 In-context Prompting Lay 0.500 0.499 0.500
0.9 In-context Prompting Lay 0.484 0.451 0.517
0.1 Fine-tuned Lay 0.507 0.495 0.520
0.5 Fine-tuned Lay 0.557 0.461 0.649
0.9 Fine-tuned Lay 0.606 0.428 0.759
Table 17: Predicted AI preference probabilities for quality. Expert readers show strong detection sensitivity in in-context prompting ($\sim$0.69 $\to$ $\sim$0.31) that reverses after fine-tuning (point estimate at 0.9 $\sim$0.68; 95% CI 0.37–0.89).
Bin Setting Reader Type Prob Prob (Low) Prob (High)
0.1 In-context Prompting Expert 0.688 0.628 0.742
0.5 In-context Prompting Expert 0.496 0.493 0.499
0.9 In-context Prompting Expert 0.306 0.254 0.363
0.1 Fine-tuned Expert 0.514 0.490 0.537
0.5 Fine-tuned Expert 0.602 0.431 0.751
0.9 Fine-tuned Expert 0.683 0.373 0.887
0.1 In-context Prompting Lay 0.454 0.421 0.488
0.5 In-context Prompting Lay 0.501 0.500 0.501
0.9 In-context Prompting Lay 0.547 0.512 0.582
0.1 Fine-tuned Lay 0.501 0.487 0.515
0.5 Fine-tuned Lay 0.509 0.403 0.614
0.9 Fine-tuned Lay 0.517 0.324 0.705
S8.3: Stylometric Correlates and Mediation
Table 18 shows that cliché density has the strongest correlation with AI detection (Pearson $r=0.60$), while readability ease is negatively correlated (r=0.23r=-0.23), suggesting AI text is marked by formulaic phrases but simpler syntax. In in-context prompting conditions, approximately 16% of the detection effect on preference is mediated through cliché density; after fine-tuning, this mediation drops to a statistically insignificant 1.3%, indicating that fine-tuning substantially reduces rather than merely masking these stylistic signatures.
Table 18: Correlations between Pangram AI detection scores and stylometric features.
Metric Pearson $r$ Spearman $\rho$ NN
readability_ease -0.232 -0.312 330
cliche_density 0.600 0.602 330
total_adjective_count 0.104 0.129 330
num_cliches 0.614 0.633 330
S8.4: Economic Analysis
Table 19 documents the cost structure for AI-based text generation. With fine-tuning costs ranging from \$22.25 to \$272.50 (median \$77.88) plus \$3 for inference to generate 100,000 words, the total AI cost represents approximately 0.3% of the \$25,000 a professional writer would charge. This $\sim$99.7% cost reduction, combined with reader preference for fine-tuned outputs, quantifies the potential economic disruption to creative writing markets.
Table 19: Cost comparison for generating 100,000 words. Fine-tuning plus inference costs (\$25–\$276) represent less than 1% of professional writer compensation (\$25,000).
NN Authors Fine-tune Min Fine-tune Med Fine-tune Max Total Min Total Med Total Max
30 \$22.25 \$77.88 \$272.50 \$25.25 \$80.88 \$275.50
Note. Inference cost of \$3 per 100k words follows fig4c_cost_summary.csv; per-token assumptions are documented in the repository.
S9. Deviations from Preregistration
This section documents deviations from the preregistered analysis plan (https://osf.io/zt4ad). All other aspects were implemented exactly as specified.
  • Model framework. The preregistration specified mixed-effects logistic regression with random intercepts for readers and prompts. We implemented logistic GLMs with CR2 cluster-robust standard errors (clustered by reader) due to convergence issues and singularity warnings in the mixed-effects models. Point estimates from both approaches were similar when convergence was achieved, and inference on the preregistered contrasts (H1, H2) remained unchanged.
  • Sample sizes (exceeded targets). We successfully recruited more participants than planned:
    • Expert readers: 28 (planned:  25)
    • Lay readers: 131 (planned: 120)
    • Fine-tuned authors: 30 (planned minimum: 10)
    All analyses used the full realized sample. No stopping rules were applied or violated; the additional data strengthens the reliability of our findings.
  • Additional analyses (not preregistered). We added:
    • Writer type × Reader type interaction tests (Type-III Wald with robust covariance) for transparency
    • Predicted probability visualizations (Figures 2C-D) to complement the odds ratio panels
    These additions provide fuller context but do not alter the conclusions drawn from the preregistered H1 and H2 contrasts.
S10. Code and Data Availability
All code and data necessary to reproduce Figures 2-4 and the associated statistical analyses will be made publicly available on paper acceptance. Please contact authors if you need it:
  • Preregistration: https://osf.io/zt4ad (Version 2, July 2025)
  • Key Analysis Scripts: The repository contains R scripts for data processing, model fitting, and figure generation. Core analyses can be reproduced by running the numbered scripts in sequence.
  • Environment: Analyses were conducted in R 4.3.1 with key packages including clubSandwich (CR2 robust SEs) and emmeans (contrasts). Full package versions are documented in the repository.
  • Data: Anonymized trial-level data in long format, with documentation of all variable definitions and transformations applied.